Calculate the pH of a solution containing 0.10 mole of acetic acid, to which is added 0.03 mole of sodium hydroxide (Ka = 1.8 x 10-5).
Answer #1

Moles of sodium acetate formed= 0.03 mol

Excess mole of acetic acid = 0.10-0.03=0.07 mol

It is acidic buffer,

pH = pKa + log(salt/acid)

= 4.74+log(0.03/0.07)

= 4.38